1.What is the probability of choosing a face card from a deck of 52 cards (face cards are jacks, queens, and kings)? What is the probability of the 2nd card being a face card if the first card was a king? (Without replacement.)
2.Compute the following and show your steps. 3! ÷ (0!*3!).
3.We have a binomial experiment with p = .8 and n = 2.
(1) Set up the probability distribution by showing all x values and their associated probabilities.
(2) Compute the mean, variance, and standard deviation

4.A class has 13 women and 18 men. If a student if chosen randomly, what is the probability the student is a woman?

Answer by stanbon(75887)

1.What is the probability of choosing a face card from a deck of 52 cards (face cards are jacks, queens, and kings)?
Ans: 12/52
--------------------
What is the probability of the 2nd card being a face card if the first card was a king? (Without replacement.)
P(fc | k) = 11/51
-----------------------

2.Compute the following and show your steps. 3! ÷ (0!*3!).
Ans:: = 3!/3! = 1
-----------------------------
3.We have a binomial experiment with p = .8 and n = 2.
(1) Set up the probability distribution by showing all x values and their associated probabilities.
P(x=0) = 0.2^2
P(x=1) = 2*0.8*0.2
-----------------------------
(2) Compute the mean, variance, and standard deviation
mean = np = 2*0.8 = 1.6
variance = npq = 1.6*0.2 = 0.32
std = sqrt(npq) = 0.5657
----------------------
4.A class has 13 women and 18 men. If a student if chosen randomly, what is the probability the student is a woman?
An: 13/31
==============================
Cheers,
Stan H.